# Dual effect of cholesterol on interfacial water dynamics in lipid membranes: Interplay between membrane packing and hydration

### v1

@@ Abstract
Water contained within biological membranes plays a critical role in maintaining the separation between intracellular and extracellular environments and facilitating biochemical processes. Variations in membrane composition and temperature lead to phase state changes in lipid membranes, which in turn influence the structure and dynamics of the associated interfacial water. In this study, molecular dynamics simulations were performed on binary membranes composed of dipalmitoylphosphatidylcholine (DPPC) or palmitoyl sphingomyelin (PSM) mixed with cholesterol (Chol). To elucidate the effects of Chol on interfacial water, we examined the orientation and hydrogen-bonding behavior of water molecules spanning from the membrane interior to the interface. As the Chol concentration increased, a transient slow down in water dynamics was observed in the gel phase at 303 K. Conversely, at higher Chol concentrations, water dynamics were accelerated relative to pure lipid membranes across all temperatures studied. Specifically, at a Chol concentration of 50%, the hydrogen bond lifetime in DPPC membranes decreased to approximately 0.5-0.7 times that of pure lipid membranes. This nonmonotonic behavior is attributed to the combined effects of membrane packing induced by Chol and a reduced density of lipid molecules in the hydrophilic region, offering key insights for modulating the dynamical properties of interfacial water.
@@ I Introduction
Cholesterol (Chol) is a representative steroid molecule present of animal cell membranes, accounting for approximately 20% to 50% of their lipid content. Its rigid ring structure and polar hydroxyl group allow Chol to integrate stably into lipid bilayers, where it promotes ordering of the hydrophobic acyl chains, leading to increased membrane condensation. [5, 6, 7, 8] These structural effects of Chol on the membrane also modulate the properties of interfacial water. [9, 10]
@@
Numerous molecular dynamics (MD) simulations have been conducted to investigate the lipid membrane–water interface. [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 7, 35, 36, 37, 38, 39, 40] MD simulation studies have also examined how Chol influences the structural organization of lipid membranes, including its effects on bilayer thickness, lipid packing, and phase behavior. [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] Specifically, clarifing the influence of Chol on the hydrogen-bond (H-bond) network within the surface hydration layer is essential for understanding membrane–water interactions at the microscopic level. Elola et al. reported that increasing the Chol concentration up to 50% enhances the diffusivity of water molecules near the membrane surface. [49] This enhancement in water mobility was attributed to the disruption of lipid-water-lipid H-bonds, while promoting the formation of water-water H-bond network. Oh et al. characterized the structure of interfacial water by modeling the HB network at lipid membrane surfaces containing Chol using a graph-based approach. [53] Their analysis indicated that increasing the Chol concentration enhances the connectivity of the H-bond network among water molecules, resulting in a structure that more closely resembles bulk-like behavior.
We previously performed MD simulations of dipalmitoylphosphatidylcholine (DPPC) and palmitoyl sphingomyelin (PSM) membranes containing 10% Chol as well as pure membranes to elucidate the dynamics of water molecules H-bonded with lipid molecules. [55] While DPPC and PSM share an identical hydrophobic tail structure, they differ in the architecture of their hydrophilic headgroups, particularly in the configuration of their polar functional groups [see Fig. 1(a) and (b)]. Notably, our MD simulations of interfacial water dynamics in two types of lipid bilayers, DPPC and PSM, revealed that water exhibits slower dynamics near PSM compared to DPPC. This behavior is primarily attributed to H-bonds formed between water hydrogen atoms and the oxygen atoms of PSM, labeled $\mathrm{O}^{\mathrm{S4}}$ and $\mathrm{O}^{\mathrm{S5}}$ [see Fig. 1(b)]. A similar retardation of interface water dynamics near PSM membranes has also been reported in previous studies. [56, 57] Furthermore, we found that the Chol slows the interfacial water dynamics, with a more pronounced effect in DPPC than in PSM. In particular, at 303 K, corresponding to the gel phase, the H-bond lifetime between lipid oxygen atoms and water hydrogen atoms increased at a Chol concentration of 10%.
Our findings and those previously reported in MD simulations indicate that Chol exhibits a dual effect of retarding and accelerating the dynamics of water molecules near the membrane interface with a significant dependence of its concentration. In this paper, we elucidate how Chol affects the H-bond dynamics of interfacial water, particularly as a function of Chol concentration.
@@ II MD simulations
The structures of the lipid molecules, DPPC, PSM, and Chol are depicted in Fig. 1. The mole fraction of Chol in the lipid membrane was set to $x_{\mathrm{Chol}}=0$, 0.1, and 0.5, with the number of water molecules maintained at a ratio of 100 per total number of lipid and cholesterol molecules (see Table 1). The CHARMM36 force field was employed for the lipid molecules (DPPC, PSM, and Chol), [58] and water was modeled using the TIP3P model optimized for compatibility with the CHARMM force field. [59] All the MD simulations were performed using GROMACS 2022.4. [60] The initial configurations of the lipid membrane systems were constructed using the CHARMM-GUI. [61, 62, 63, 64] System equilibration was carried out following the CHARMM-GUI protocol, as described in our previous study. [55] Three initial configurations were prepared for each $x_{\mathrm{Chol}}$. The temperatures investigated were 303 K and 323 K. The production runs of 10 ns in the $NPT$ ensemble. Each production run was performed three times per configuration with different initial atomic velocities assigned according to the Maxwell–Boltzmann distribution, resulting in a total of nine trajectories for each combination of $x_{\mathrm{Chol}}$ and temperature.
@@ III.1 Area per lipid
Figure 2 presents snapshots of DPPC-Chol systems at 303 K, illustrating the effect of varying Chol concentration, $x_{\mathrm{Chol}}$. In this study, the total number of lipid molecules was fixed, as described in Table 1. Thus, increasing $x_{\mathrm{Chol}}$ led to a corresponding increase in the simulation box size. It is also observed that $x_{\mathrm{Chol}}$ influences the bending behavior of membranes and the orientation of the hydrophobic group of the lipid molecules. The top-down snapshots in Fig. 2 reveal that at $x_{\mathrm{Chol}}=0.1$, Chol is stably incorporated into the membrane and exhibits strong lateral aggregation, with minimal exposure at the membrane interface. In contrast, at $x_{\mathrm{Chol}}=0.5$, the presence of Chol becomes apparent at the membrane interface, which is regarded as the effective dilution of lipid molecules compared to the system with $x_{\mathrm{Chol}}=0.1$.
@@ III.2 Water distribution near membrane interface
Figure 4(a) and (b) show the ratio $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ at 303 K for DPPC and PSM, respectively. The results of $x_{\mathrm{Chol}}=0$ and 0.1 are consistent with those reported in our previous study. [55] Figure S1 of the supplementary material shows the results at 323 K. Here, the $z^{\prime}$ axis is divided into three distinct regions: the membrane interior (Region 1), the interface (Region 2), and the bulk aqueous phase (Region 3). [31, 33] We defined Region 1 as $z^{\prime}<-0.25$ nm, Region 2 as $-0.25~\mathrm{nm}<z^{\prime}<0.85\mathrm{nm}$, and Region 3 as $z^{\prime}>0.85~\mathrm{nm}$, as indicated by the dashed vertical lines in Fig. 4.
@@
Both DPPC and PSM exhibit a non-monotonic dependence on $x_{\mathrm{Chol}}$ in Region 2 at 303 K. In particular, for DPPC, the number density of water at the membrane interface decreased from $x_{\mathrm{Chol}}=0$ to $x_{\mathrm{Chol}}=0.1$, but increases markedly at $x_{\mathrm{Chol}}=0.5$. This increase in $\rho(z^{\prime})$ is consistent with the results performed by Elola et al. [49] As illustrated in Fig. S1 of the supplementary material, at 323 K, the number density of water in Region 2 shows little variation between $x_{\mathrm{Chol}}=0$ and $x_{\mathrm{Chol}}=0.1$. However, at $x_{\mathrm{Chol}}=0.5$, it increases noticeably, similar to the trend observed at 303 K. This non-monotonic behavior of $\rho(z^{\prime})$ appears to contrast with the results in Fig. 3, which indicate that increasing $x_{\mathrm{Chol}}$ leads to a more condensed lipid membrane state. This apparent discrepancy can be interpreted as follows: Chol, being shorter than typical lipid molecules, resides near the membrane center. While its hydroxyl group and the lipid phosphate groups are in close proximity, they do not extend fully to the membrane interface, particularly at a low Chol concentration. However, at high Chol concentrations, the lateral spacing between lipid headgroups increases, leading to greater exposure of Chol at the interface, which can be observed in the snapshot in Fig. 2. This increased exposure facilitates the accumulation of water molecules at the interface.
@@
Compared to DPPC, PSM exhibits small changes in $\rho(z^{\prime})$ at $x_{\mathrm{Chol}}=0.1$ from $x_{\mathrm{Chol}}=0$, while at $x_{\mathrm{Chol}}=0.5$, it shows a compatible increase in the water number density at Region 2, similar to the behavior observed in DPPC. As mentioned in Sec. III.1, the area per lipid is smaller for PSM than for DPPC in the absence of Chol (see Fig. 3), indicating that PSM forms a more tightly packed membrane. Consequently, the exclusion effect of Chol on water molecules is less pronounced in PSM. Eventually, as the Chol concentration increases, the lateral spacing between lipid headgroups expands, promoting the redistribution of water molecules from the membrane interior toward the interface.
We next examined the H-bond number, $N_{\mathrm{HB}}$, formed by water molecules, as a function of $z^{\prime}$, as is shown in Fig. 4 (c) and (d), for DPPC and PSM, respectively. The corresponding results at 323 K are displayed in Fig. S1 (c) and (d) of the supplementary material. An H-bond was identified based on the geometric criteria: an oxygen-oxygen distance $r_{\mathrm{oo}}<0.35~\mathrm{nm}$ and a hydrogen-donor-acceptor angle $0^{\circ}<\beta<30^{\circ}$[68] Note that when counting $N_{\mathrm{HB}}$, we included not only those between water molecules but also those formed between water molecules and lipid oxygen atoms, ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D6}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, as well as Chol oxygen ($\mathrm{O}^{\mathrm{Chol}}$) (see Fig. 1). The value of $N_{\mathrm{HB}}$ decreases progressively towards the membrane interior in both DPPC and PSM. In Region 1, the presence of Chol noticeably reduces $N_{\mathrm{HB}}$, compared to the Chol-free system.
@@
Figure 5: Two-dimensional plots of joint distribution of $\rho(z^{\prime},\cos\theta)$, normalized by $\rho(z^{\prime})$, at 303K for $x_{\mathrm{Chol}}=0$ [(a) and (d)], 0.1 [(b) and (e)], and 0.5 [(c) and (f)]. Panels (a), (b), and (c) correspond to DPPC, while panels (d), (e), and (f) corresponds to PSM.
@@
In addition, we analyzed the normalized number of H-bonds formed by oxygen atoms surrounding the water molecules at each position $z^{\prime}$. Specifically, the number of oxygen atoms located within $r_{\mathrm{OO}}<0.35$ nm is denoted as $N_{\mathrm{NN}}$, representing the number of nearest neighbors. The normalized number of H-bonds, $N_{\mathrm{HB}}/N_{\mathrm{NN}}$, was then evaluated as a function of $z^{\prime}$, with the results shown in Fig. 4(e) and (f). Note that $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ denotes the average fraction of neighboring oxygen atoms that form H-bonds at each position $z^{\prime}$[69, 70] In the bulk region (Region 3), the value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ is approximately 0.6, indicating that about 60% of the oxygen atoms in the nearest neighbors participate in H-bonding. The value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ increases toward the interior of the membrane (Region 1) in the Chol-free system. In other words, when oxygen atoms are present nearby, H-bonds are formed with a higher probability than in the bulk, primarily due to the proximity of oxygen atoms from lipid molecules. However, at $x_{\mathrm{Chol}}=0.5$, the Chol suppressed H-bond formation in the membrane interior for both DPPC and PSM. This suppression can be attributed to the reduced number of H-bonds in that region caused by Chol, as shown in Fig. 4(c) and (d).

### v2

@@ Abstract
Water contained within biological membranes plays a critical role in maintaining the separation between intracellular and extracellular environments and facilitating biochemical processes. Variations in membrane composition and temperature lead to phase state changes in lipid membranes, which in turn influence the structure and dynamics of the associated interfacial water. In this study, molecular dynamics simulations were performed on binary membranes composed of dipalmitoylphosphatidylcholine (DPPC) or palmitoyl sphingomyelin (PSM) mixed with cholesterol (Chol). To elucidate the effects of Chol on interfacial water, we examined the orientation and hydrogen-bonding behavior of water molecules spanning from the membrane interior to the interface. As the Chol concentration increased, a transient slow down in water dynamics was observed in the ripple phase at 303 K. Conversely, at higher Chol concentrations, water dynamics were accelerated relative to pure lipid membranes across all temperatures studied. Specifically, at a Chol concentration of 50%, the hydrogen bond lifetime in DPPC membranes decreased to approximately 0.5-0.7 times that of pure lipid membranes. This nonmonotonic behavior is attributed to the combined effects of membrane packing induced by Chol and a reduced density of lipid molecules in the hydrophilic region, offering key insights for modulating the dynamical properties of interfacial water.
@@ I Introduction
Cholesterol (Chol) is a representative steroid molecule present of animal cell membranes, accounting for approximately 20% to 50% of their lipid content. For instance, the membranes of human blood cells contain palmitoyl sphingomyelin (PSM) and analogs of dipalmitoylphosphatidylcholine (DPPC), and their Chol content is approximately 50%. [5] Owing to its rigid ring structure and polar hydroxyl group, Chol integrates stably into lipid bilayers, where it promotes ordering of the hydrophobic acyl chains, leading to increased membrane condensation. [6, 7, 8, 9] These structural effects of Chol on the membrane also modulate the properties of interfacial water. [10, 11]
@@
Numerous molecular dynamics (MD) simulations have been conducted to investigate the lipid membrane–water interface. [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 8, 36, 37, 38, 39, 40, 41] MD simulation studies have also examined how Chol influences the structural organization of lipid membranes, including its effects on bilayer thickness, lipid packing, and phase behavior. [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] Specifically, clarifing the influence of Chol on the hydrogen-bond (H-bond) network within the surface hydration layer is essential for understanding membrane–water interactions at the microscopic level. Elola and Rodriguez reported that increasing the Chol concentration of 30% and 50% enhances the diffusivity of water molecules near the membrane surface. [50] This enhancement in water mobility was attributed to the disruption of lipid-water-lipid H-bonds, while promoting the formation of water-water H-bond network. Oh et al. characterized the structure of interfacial water by modeling the HB network at lipid membrane surfaces containing Chol using a graph-based approach. [54] Their analysis indicated that increasing the Chol concentration enhances the connectivity of the H-bond network among water molecules, resulting in a structure that more closely resembles bulk-like behavior.
We previously performed MD simulations of dipalmitoylphosphatidylcholine (DPPC) and palmitoyl sphingomyelin (PSM) membranes containing 10% Chol as well as pure membranes to elucidate the dynamics of water molecules H-bonded with lipid molecules. [56] While DPPC and PSM share an identical hydrophobic tail structure, they differ in the architecture of their hydrophilic headgroups, particularly in the configuration of their polar functional groups [see Fig. 1(a) and (b)]. Notably, our MD simulations of interfacial water dynamics in two types of lipid bilayers, DPPC and PSM, revealed that water exhibits slower dynamics near PSM compared to DPPC. This behavior is primarily attributed to H-bonds formed between water hydrogen atoms and the oxygen atoms of PSM, labeled $\mathrm{O}^{\mathrm{S4}}$ and $\mathrm{O}^{\mathrm{S5}}$ [see Fig. 1(b)]. A similar retardation of interface water dynamics near PSM membranes has also been reported in previous studies. [57, 58] Furthermore, we found that the Chol slows the interfacial water dynamics, with a more pronounced effect in DPPC than in PSM. Our findings and those previously reported in MD simulations indicate that Chol exhibits a dual effect of retarding and accelerating the dynamics of water molecules near the membrane interface with a significant dependence of its concentration. In this paper, we elucidate how Chol affects the H-bond dynamics of interfacial water, particularly as a function of Chol concentration.
@@ II MD simulations
[…]
The MD setting are the same as that in our previous study. The time step was set to 2 fs. The system temperature was controlled using the Nosé–Hoover thermostat with a coupling constant of 1 ps. [68, 69] We employed Parrinello–Rahman barostat with a coupling constant of 5 ps and a compressibility of $4.5\times 10^{-5}~\text{bar}^{-1}$[70] Long-range electrostatic interactions were treated using the smooth Particle-Mesh Ewald (SPME) method with a 1.2 nm cut-off. [71] Short-range van der Waals interactions were handled with a force-switch modifier between 1.0 and 1.2 nm. Neighbor lists were updated every 20 steps using the Verlet cutoff scheme. All bonds involving hydrogen atoms were constrained using the LINCS algorithm to allow the 2 fs time step. [72]
@@ III.1 Area per lipid
Figure 2 presents snapshots of DPPC-Chol systems at 303 K, illustrating the effect of varying Chol concentration, $x_{\mathrm{Chol}}$. In this study, the total number of lipid molecules was fixed, as described in Table 1. Thus, increasing $x_{\mathrm{Chol}}$ led to a corresponding increase in the simulation box size. As shown in Fig. 2(a), the pure DPPC membrane at 303 K exhibits the periodic modulation of the bilayer surface, a characteristic feature of ripple phase. The phase behavior we obtained is consistent with the experimental findings, as 303 K is close to the temperature range of 308–314 K, in which the ripple phase is experimentally observed. [74, 75] The top-down snapshots in Fig. 2 reveal that at $x_{\mathrm{Chol}}=0.1$, Chol is stably incorporated into the membrane and exhibits strong lateral aggregation, with minimal exposure at the membrane interface. In contrast, at $x_{\mathrm{Chol}}=0.5$, the presence of Chol becomes apparent at the membrane interface, which is regarded as the effective dilution of lipid molecules compared to the system with $x_{\mathrm{Chol}}=0.1$. From the radial distribution functions (RDFs) on the $xy$-plane between the phosphorus atoms and between phosphorus and cholesterol oxygen ($\mathrm{O}^{\mathrm{Chol}}$) atoms, we confirm that no macroscopic phase separation is observed and that the lipids (DPPC and PSM) and cholesterol are miscible at $x_{\mathrm{Chol}}=0$, 0.1, and 0.5, as the RDFs converge to unity in the long-range limit for all cases (see Figs. S2 and S3 of the supplementary material).
@@ III.2 Water distribution near membrane interface
Figure 4(a) and (b) show the ratio $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ at 303 K for DPPC and PSM, respectively. The results of $x_{\mathrm{Chol}}=0$ and 0.1 are consistent with those reported in our previous study. [56] Figure S4(a) and (b) of the supplementary material shows the results at 323 K. Here, the $z^{\prime}$ axis is divided into three distinct regions: the membrane interior (Region 1), the interface (Region 2), and the bulk aqueous phase (Region 3). [32, 34] We defined Region 1 as $z^{\prime}<-0.25$ nm, Region 2 as $0.25~\mathrm{nm}<z^{\prime}<0.85\mathrm{nm}$, and Region 3 as $z^{\prime}>0.85~\mathrm{nm}$, as indicated by the dashed vertical lines in Fig. 4.
@@
Both DPPC and PSM exhibit a non-monotonic dependence on $x_{\mathrm{Chol}}$ in Region 2 at 303 K. In particular, for DPPC, the number density of water at the membrane interface decreases from $x_{\mathrm{Chol}}=0$ to $x_{\mathrm{Chol}}=0.1$, but increases markedly at $x_{\mathrm{Chol}}=0.5$. This increase in $\rho(z^{\prime})$ is consistent with the results performed by Elola and Rodriguez [50] As illustrated in Fig. S4 of the supplementary material, at 323 K, the number density of water in Region 2 shows little variation between $x_{\mathrm{Chol}}=0$ and $x_{\mathrm{Chol}}=0.1$. However, at $x_{\mathrm{Chol}}=0.5$, it increases noticeably, similar to the trend observed at 303 K. This non-monotonic behavior of $\rho(z^{\prime})$ appears to contrast with the results in Fig. 3, which indicate that increasing $x_{\mathrm{Chol}}$ leads to a more condensed lipid membrane state. This apparent discrepancy can be interpreted as follows: Chol, being shorter than typical lipid molecules, resides near the membrane center. While its hydroxyl group and the lipid phosphate groups are in close proximity, they do not extend fully to the membrane interface, particularly at a low Chol concentration. However, at high Chol concentrations, the lateral spacing between lipid headgroups increases, leading to greater exposure of Chol at the interface, which can be observed in the snapshot in Fig. 2. This increased exposure facilitates the accumulation of water molecules at the interface.
@@
Compared to DPPC, PSM exhibits small changes in $\rho(z^{\prime})$ at $x_{\mathrm{Chol}}=0.1$ from $x_{\mathrm{Chol}}=0$, while at $x_{\mathrm{Chol}}=0.5$, it shows a compatible increase in the water number density at Region 2, similar to the behavior observed in DPPC. As mentioned in Sec. III.1, the area per lipid is smaller for PSM than for DPPC in the absence of Chol (see Fig. 3), indicating that PSM forms a more tightly packed membrane. Consequently, the exclusion effect of Chol on water molecules is less pronounced in PSM. As the Chol concentration increases up to $x_{\mathrm{Chol}}=0.5$, the lateral spacing between lipid headgroups expands, leading to a shift of the rising edge of the water density profile toward the interfacial region.
[…]
@@
Figure 5: Left panel: Schematic illustration of water orientation. Right panels: Two-dimensional plots of joint distribution of $\rho(z^{\prime},\cos\theta)$, normalized by $\rho(z^{\prime})$, at 303K for $x_{\mathrm{Chol}}=0$ [(a) and (d)], 0.1 [(b) and (e)], and 0.5 [(c) and (f)]. Panels (a), (b), and (c) correspond to DPPC, while panels (d), (e), and (f) corresponds to PSM.
@@
In addition, we analyzed the normalized number of H-bonds formed by oxygen atoms surrounding the water molecules at each position $z^{\prime}$. Specifically, the number of oxygen atoms located within $r_{\mathrm{OO}}<0.35$ nm is denoted as $N_{\mathrm{NN}}$, representing the number of nearest neighbors. The normalized number of H-bonds, $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$, was then evaluated as a function of $z^{\prime}$, with the results shown in Fig. 4(e) and (f). The corresponding results at 323 K are displayed in Fig. S4 (e) and (f) of the supplementary material. Note that $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ denotes the average fraction of neighboring oxygen atoms that form H-bonds at each position $z^{\prime}$[82, 83] In the bulk region (Region 3), the value of $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ is approximately 0.6, indicating that about 60% of the oxygen atoms in the nearest neighbors participate in H-bonding. The value of $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ increases toward the interior of the membrane (Region 1) in the Chol-free system. In other words, when oxygen atoms are present nearby, H-bonds are formed with a higher probability than in the bulk, primarily due to the proximity of oxygen atoms from lipid molecules. However, at $x_{\mathrm{Chol}}=0.5$, $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ is reduced in the membrane interior, even though $\Braket{N_{\mathrm{HB}}}$ is hardly changed. Since $\Braket{N_{\mathrm{NN}}}$ is higher at $x_{\mathrm{Chol}}=0.5$ than at $x_{\mathrm{Chol}}=0.1$ (see Fig. S5 of the supplementary material), the decrease in $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ indicates that increasing Chol to $x_{\mathrm{Chol}}=0.5$ raises the number of neighboring oxygen atoms that do not participate in H-bonding.